# Wave dynamics governing vortex breakdown in smooth Euler flows

### v1

@@ Abstract.
We consider the three-dimensional incompressible Euler equations in a setting where a vortex is transported by a prescribed Lagrangian flow. We show that vortex breakdown is governed by wave dynamics generated by the underlying transport flow. The key idea is to avoid any singular integral representation of the pressure term and instead construct an effective Lagrangian coordinate system in which the associated Lie bracket vanishes identically.
@@
Euler equations, vortex breakdown, Frenet-Serret formulas
@@ 1. Introduction
One of the most important features of Navier–Stokes turbulence is that it is not random but is instead organized around vortex stretching dynamics. More precisely, recent direct numerical simulations of Navier–Stokes turbulence at sufficiently high Reynolds numbers [2, 3, 10, 11] have reported the existence of a hierarchical structure of vortex-stretching motions. In particular, Goto, Saito, and Kawahara [3] clearly observed that turbulence at sufficiently high Reynolds numbers in a periodic cube exhibits a self-similar hierarchy of antiparallel vortex-tube pairs, which is sustained by the generation of smaller-scale vortices through stretching within larger-scale strain fields. This observation was further investigated by Y–Goto–Tsuruhashi [14] and Tsuruhashi–Goto–Oka–Y [13] (see also [5, 6, 7] for related mathematical results). Following this line of research, attention naturally turns to the dynamics of vortex stretching and the mechanisms underlying the subsequent breakdown of vortices.
With this perspective in mind, in this paper we provide a mathematical analysis of vortex breakdown in a scenario where a vortex is transported by a prescribed Lagrangian flow. Let $\Phi$ denote the Lagrangian flow map associated with the incompressible Euler equations, written in Lagrangian form (omitting the explicit volume-preserving constraint) as follows:
@@
and we interpret $Z$ as describing the time evolution of swirling particles around the vortex axis. The initial data $\alpha_{1}(0),\alpha_{2}(0)\in\mathbb{R}$ can be chosen arbitrarily.
Remark 1.
The time derivatives $\partial_{t}\alpha_{1}(0)$ and $\partial_{t}\alpha_{2}(0)$ also serve as initial data. In this case, we obtain the relations
$\begin{split}\nabla_{x}\bigl(\partial_{t}\Phi(0,x)\cdot\tau(0)\bigr)\big|_{x=\ell(s)}\cdot n&=\partial_{t}\alpha_{1}(0)-\kappa(0)v(0),\\ \nabla_{x}\bigl(\partial_{t}\Phi(0,x)\cdot\tau(0)\bigr)\big|_{x=\ell(s)}\cdot b&=\partial_{t}\alpha_{2}(0).\end{split}$
These formulas show that $\partial_{t}\alpha_{1}(0)$ and $\partial_{t}\alpha_{2}(0)$ are determined by the Eulerian shear of the background initial velocity field and the curvature of the particle trajectory. If the Eulerian shear vanishes at the initial time, then
$\partial_{t}\alpha_{1}(0)=\kappa(0)v(0),\qquad\partial_{t}\alpha_{2}(0)=0,$
and the subsequent transverse dynamics are driven purely by the Lagrangian geometry of the vortex curve. This shear-free setting is physically natural and theoretically significant: while strong shear can always trigger vortex instabilities, the absence of shear isolates the intrinsic geometric mechanism of vortex breakdown.
Remark 2.
We now give a definition of the occurrence of vortex breakdown at a time $t_{0}$. Define the axis of swirling particles $\zeta(t)$ by
(6) $\zeta(t):=\tau(t)-\alpha_{1}(t)\,n(t)-\alpha_{2}(t)\,b(t).$
On the other hand, the vortex axis $\zeta^{*}(t)$ is defined by
$\zeta^{*}(t):=\partial_{s}\Phi(t,s,R_{1},R_{2})\big|_{R_{1}=R_{2}=0}.$
We say that vortex breakdown occurs at time $t_{0}$ if
$\cos^{-1}\!\left(\frac{\zeta^{*}(t)}{|\zeta^{*}(t)|}\cdot\frac{\zeta(t)}{|\zeta(t)|}\right)\longrightarrow\frac{\pi}{2}\qquad(t\to t_{0}).$
Then the main theorem is as follows:
@@
The quantities $\alpha_{1}$ and $\alpha_{2}$, which govern the occurrence of vortex breakdown, satisfy the following second–order wave equations:

### v2

# Nonlinear Scale-Local Deformations of Vortex Rings in Smooth Euler Flows

@@ Abstract.
We consider the incompressible three-dimensional Euler equations in a setting where a vortex ring undergoes radially expanding Lagrangian transport. To clarify the fundamental mechanisms responsible for nonlinear scale-local deformations of the vortex structure, we develop a geometric Lagrangian framework that avoids singular integral representations of the pressure and leads to the derivation of a novel wave equation governing the axis of swirling particles. Within this framework, we identify purely nonlinear mechanisms responsible for the emergence of scale-local deformations of the vortex structure, with the aid of a machine-learning-based analysis.
@@
Euler equations, Frenet-Serret formulas, Optuna
@@ 1. Introduction
One of the most important features of Navier–Stokes turbulence is that it is not random but is instead organized around vortex-stretching dynamics. More precisely, recent direct numerical simulations of Navier–Stokes turbulence at sufficiently high Reynolds numbers [2, 3, 11, 12] have reported the existence of a hierarchical structure of vortex-stretching motions. In particular, Goto-Saito-Kawahara [3] clearly observed that turbulence at sufficiently high Reynolds numbers in a periodic cube exhibits a self-similar hierarchy of antiparallel vortex-tube pairs, which is sustained by the generation of smaller-scale vortices through stretching within larger-scale strain fields. This observation was further investigated by Y–Goto–Tsuruhashi [15] and Tsuruhashi–Goto–Oka–Y [14] (see also [5, 6, 7] for related mathematical results). Following this line of research, attention naturally turns to the scale-local deformations of vortical structures, which play a central role in the energy cascade.
A natural setting in which such scale-local deformation mechanisms can be examined in a controlled manner is provided by the collision of vortex rings. Many studies have investigated vortex-ring collisions (see [10] and references therein), and the main findings can be summarized as follows. When two vortex rings of equal strength and opposite circulation collide head-on, a strong strain field is generated. It is well known that the dominant instability mechanism depends on the Reynolds number. At low Reynolds numbers, long-wavelength Crow-type instabilities prevail, leading to relatively mild deformations. In contrast, at high Reynolds numbers, short-wavelength perturbations at the scale of the vortex core radius become dominant, and the dynamics are governed by the elliptical instability. However, the elliptical instability is derived from a linear stability analysis and does not, by itself, provide a rigorous description of the nonlinear evolution.
With this perspective in mind, in this paper, we develop a geometric Lagrangian framework for the three-dimensional incompressible Euler equations that avoids singular integral representations of the pressure. Within this framework, we rigorously derive a novel wave equation governing the evolution of the axis of swirling particles, and we further employ a machine-learning-based analysis to identify purely nonlinear mechanisms responsible for the emergence of scale-local deformations of the vortex structure.
2. Derivation of a Novel Wave Equation
Let $\Phi$ denote the Lagrangian flow map associated with the incompressible Euler equations, written in Lagrangian form (omitting the explicit volume-preserving constraint) as follows:
@@ 2. Derivation of a Novel Wave Equation
and we interpret $Z$ as describing the time evolution of swirling particles around the vortex axis. We are now in a position to state the following key theorem.
@@
The quantities $\alpha_{1}$ and $\alpha_{2}$ satisfy the following second–order wave equations:
@@ 3. Proof of Theorem 1
4. Numerical Implementation
In this section, we present a numerical implementation to investigate how the alignment between the vortex axis and the axis of swirling particles is maintained during the time evolution of a vortex ring. In particular, we demonstrate that such alignment cannot be preserved by pure radial expansion alone, and that a nontrivial deformation of the ring plays an essential role. We first prescribe an accelerating transport flow that expands in the radial direction, characterized by a scalar function $\Gamma(t)$ satisfying
$\Gamma(t_{0})=0,\qquad\Gamma(t)>0,\qquad\Gamma^{\prime}(t)>0,\qquad\Gamma^{\prime\prime}(t)>0\quad\text{for}\quad t\in(t_{0},t_{1}).$
The time evolution of the ring shape, for $t\in[t_{0},t_{1})$ and $s\in[0,1)$, is defined by
$\Phi(t,s)=\Bigl(R(s,\delta)+\Gamma(t)+\gamma_{1}(t,s)\Bigr)\begin{pmatrix}\cos 2\pi s\\ \sin 2\pi s\\ 0\end{pmatrix}+\gamma_{2}(t,s)\begin{pmatrix}0\\ 0\\ 1\end{pmatrix}$.
Here,
$R(s,\delta):=\frac{(1+\delta)(1-\delta)}{2\sqrt{\bigl((1-\delta)\cos s\bigr)^{2}+\bigl((1+\delta)\sin s\bigr)^{2}}}$
represents the initial ring shape with elliptic radius, where $\delta>0$ is a small ellipticity parameter. The functions $\gamma_{1}(t,s)$ and $\gamma_{2}(t,s)$ describe a time-dependent wavy deformation and are given by
$\begin{split}\gamma_{1}(t,s)&=\frac{1}{K+1}\sum_{j=0}^{J}\sum_{k=0}^{K}\Bigl(c^{11}_{jk}(t-t_{0})^{j+1}\sin 2\pi ks+c^{12}_{jk}(t-t_{0})^{j+1}\cos 2\pi ks\Bigr),\\ \gamma_{2}(t,s)&=\frac{1}{K+1}\sum_{j=0}^{J}\sum_{k=0}^{K}\Bigl(c^{21}_{jk}(t-t_{0})^{j+1}\sin 2\pi ks+c^{22}_{jk}(t-t_{0})^{j+1}\cos 2\pi ks\Bigr),\end{split}$
where the coefficients $\{c^{\ell m}_{jk}\}_{jk\ell m}\subset\mathbb{R}$ are treated as learnable parameters. We now give a definition of the axis of swirling particles $\zeta(t,s)$ by
(10) $\zeta(t,s):=\tau(t,s)-\alpha_{1}(t,s)n(t,s)-\alpha_{2}(t,s)b(t,s).$
On the other hand, the vortex axis $\zeta^{*}(t,s)$ is defined by
$\zeta^{*}(t,s):=\partial_{s}\Phi(t,s,R_{1},R_{2})\big|_{R_{1}=R_{2}=0}.$
Recall taht $\alpha_{1}$ and $\alpha_{2}$ satisfy the following wave-type equations:
$\alpha_{1}^{\prime\prime}=\frac{v^{\prime\prime}}{v}\alpha_{1}+2v\kappa^{\prime}+4v^{\prime}\kappa,\qquad\alpha_{2}^{\prime\prime}=\frac{v^{\prime\prime}}{v}\alpha_{2},$
where $v(t,s):=|\partial_{t}\Phi(t,s)|$ and $\kappa(t,s)$ denotes the curvature of the curve $\cup_{t}\Phi(t,s)$. For the machine learning implementation, we choose the initial data $\alpha_{1}(0,s)$, $\alpha_{2}(0,s)$, $\partial_{t}\alpha_{1}(0,s)$, and $\partial_{t}\alpha_{2}(0,s)$ so that
$\frac{\zeta^{*}(0,s)}{|\zeta^{*}(0,s)|}\cdot\frac{\zeta(0,s)}{|\zeta(0,s)|}=1,\qquad\partial_{t}\left(\frac{\zeta^{*}(t,s)}{|\zeta^{*}(t,s)|}\cdot\frac{\zeta(t,s)}{|\zeta(t,s)|}\right)\bigg|_{t=0}=0\quad\text{for}\quad s\in[0,1).$
These conditions ensure that the vortex axis and the axis of swirling particles are perfectly aligned at the initial time. If the two axes are not parallel, the resulting rotating flow tends to exhibit shear. From a physical viewpoint, it is therefore natural to assume that the system favors configurations in which these axes remain aligned during the evolution, thereby minimizing shear. Motivated by this consideration, we introduce the following mean absolute directional correlation (MADC) as a functional:
$\mathrm{MADC}:=\frac{1}{t_{1}-t_{0}}\int_{t_{0}}^{t_{1}}\int_{0}^{1}\left|\frac{\zeta^{*}(t,s)}{|\zeta^{*}(t,s)|}\cdot\frac{\zeta(t,s)}{|\zeta(t,s)|}\right|\,ds\,dt,$
which serves as an objective functional to be maximized. In the numerical implementation, we set
$\Gamma(t)=1-\cos(12\pi t),\qquad(t_{0}=1/48,\;t_{1}=1/24),$
and adopt the following hyperparameters:
  • $\delta=0.02$ (ellipticity parameter),
  • $J=20$ (number of polynomial terms),
  • $K=10$ (number of Fourier modes).
The time interval $t_{0},t_{1}]$ is discretized into $32$ steps, and the angular parameter interval $[0,1)$ is discretized into $128$ points. For optimization, we employ Bayesian optimization using the Optuna library:
  • First 10,00010{,}000 trials: quasi–Monte Carlo sampling,
  • Subsequent $50$ trials: Bayesian optimization,
  • Parameter bounds: $|c^{\ell m}_{jk}|\leq 30.0$.
The numerical results are shown in Figures 1 and 2.
Figure 1. Initial vortex rings (the same configuration). Left: Bayesian optimization, Right: $\{c^{\ell m}_{jk}\}_{jk\ell m}=0$
Figure 2. Vortex rings at the terminal time. Left: Bayesian optimization, Right: $\{c^{\ell m}_{jk}\}_{jk\ell m}=0$
Qualitatively, the results indicate that maintaining alignment between the vortex axis and the axis of swirling particles throughout the evolution requires the presence of a nontrivial wavy deformation. Conversely, larger values of MADC are obtained only when such wavy deformations are included, indicating that these deformations play a fundamental role in generating the discrepancies captured by the error metric. At the present stage, our analysis remains qualitative. Nevertheless, despite prescribing ten Fourier modes in the Bayesian learning procedure, the representations inferred from the image data consistently indicate that the effective dynamics are dominated by five to six Fourier modes. This observation provides strong evidence that the deformation of the vortex ring is governed by scale-local mechanisms. A central direction for future work is to undertake a detailed qualitative investigation of this scale locality.